# Large Continual Instruction Assistant

### v3

@@ Abstract
Continual Instruction Tuning (CIT) is adopted to continually instruct Large Models to follow human intent data by data. It is observed that existing gradient update would heavily destroy the performance on previous datasets during CIT process. Instead, Exponential Moving Average (EMA), owns the ability to trace previous parameters, which can aid in decreasing forgetting. Nonetheless, its stable balance weight fails to deal with the ever-changing datasets, leading to the out-of-balance between plasticity and stability. In this paper, we propose a general continual instruction tuning framework to address the challenge. Starting from the trade-off prerequisite and EMA update, we propose the plasticity and stability ideal condition. Based on Taylor expansion in the loss function, we find the optimal balance weight can be automatically determined by the gradients and learned parameters. Therefore, we propose a stable-plasticity balanced coefficient to avoid knowledge confusion. Based on the semantic similarity of the instructions, we can determine whether to retrain or expand the training parameters and allocate the most suitable parameters for the testing instances. Extensive experiments across multiple continual instruction tuning benchmarks demonstrate that our approach not only enhances anti-forgetting capabilities but also significantly improves overall continual tuning performance. For example, based on LLaVA-7B, the forgetting is reduced from 5.42 to 1.93. Our code will be made publicly available soon.
@@ 1 Introduction
Large Foundation Modals (LFMs) have demonstrated remarkable capabilities in multi-task understanding and generation (Li et al., 2023; Zhu et al., 2023; Achiam et al., 2023; Touvron et al., 2023a). It generally exits two stages: large scale pre-training and instruction-tuning. Instruction tuning is extremely significant due to guiding LFMs (Chen et al., 2024b; Zheng et al., 2023; Touvron et al., 2023b) in following human intent and aligning different modalities (Li et al., 2024; Liu et al., 2023), which seems to be an essential technique to enhance the capabilities and controllability of LFMs e.g. LLMs, MLLMs.
As knowledge continuously evolves with the development of human society, new instructions are constantly generated, e.g. the emergence of new concepts or disciplines. How to enable existing LFMs to assimilate novel instructions and undergo self-evolution becomes the key challenge (Zhu et al., 2024). To accommodate the new instructions, the most effective strategy is incorporating both old and new instructions for joint training. However, even such relatively lightweight fine-tuning is unaffordable. Furthermore, directly fine-tuning these new instructions would destroy the pre-training knowledge, e.g. catastrophic forgetting (Goodfellow et al., 2013; Li et al., 2019; Nguyen et al., 2019).
@@
Continual Instruction Tuning (CIT) is proposed to address this challenge (Zhang et al., 2023; He et al., 2023; Chen et al., 2024a). Taking MLLMs as a case study, previous methods, EProj (He et al., 2023), FwT(Zheng et al., 2024), and CoIN (Chen et al., 2024a), utilize the model expansion framework by continually adding new branches for the novel instructions, which therefore has less impact on the old knowledge. However, they suffer from memory explosion and high computational cost problems. On the other hand, continually full fine-tuning (FFT) downstream datasets with single branch architecture would destroy the pre-trained parameters, and greatly reduce the zero-shot generalization performance of MLLMs (Zhai et al., 2023).
@@
In order to validate our method, for MLLMs, we adopt LLaVA-1.5 (Liu et al., 2024a) and Qwen-VL (qwen) as the backbone and insert the efficient-tuning parameters: LoRA (Hu et al., 2022) in the LLM. We continually fine-tune on multimodal instruction datasets and verify performances of the tuned model. Furthermore, we also implement our method with the LM-adapted version of T5-small on the NLP continual instruction tuning tasks (t5; Zhang et al., 2023). Experimental results consistently demonstrate excellent anti-forgetting, and continual tuning performance of our method. In summary, the contributions of this paper are as follows:
@@
Generalized Application and Limited Tuning Costs. Our method is model-free and can be easily applied to a wide range of CIT methods. Additionally, it costs limited tuning resources, due to instruction grouping strategy.
State-of-The-Art Performance. To the best of our knowledge, our method shows the best comprehensive continual tuning performance, especially significant enhancements in the performance of the baseline (as shown in Figure LABEL:radar).
@@ 2 Related Work
Large Foundation Models. LFMs own strong reasoning and complex contextual understanding capabilities, which can generate text sequences according to the input images and texts. With a frozen large language model and visual tower, BLIP2 (Li et al., 2023) bridged the gap between image and text modality through Qformer structure. Inspired by the notable reasoning power of instruction tuning in LLMs, e.g., LLaMA (Touvron et al., 2023a) and GPT (Floridi & Chiriatti, 2020; Achiam et al., 2023), LLaVA (Liu et al., 2024b) and MiniGPT4 (Zhu et al., 2023) adopted instruction tuning with only training a linear projection layer to align the image-text modalities. Recently, LLaVA-1.5 (Liu et al., 2024a) and MiniGPT5 (Zheng et al., 2023) refined the instructions and improved the performance across a wider range of multi-modality datasets.
Continual Instruction Tuning Work: (Zhang et al., 2023) first proposed the continual instruction tuning with LLMs. Following it, (He et al., 2023) and (Chen et al., 2024a) designed distinct cross-modality continual instruction tuning benchmarks with InstructBLIP and LLaVA, respectively. Specifically, (He et al., 2023) adopted unique linear projection layers for each tuning dataset and utilized a key-query driven mechanism to infer the dataset identifier. (Chen et al., 2024a) employed the MOELoRA paradigm, assigning different LoRA weights to different datasets based on expert knowledge. Nevertheless, they still faced the issue of forgetting due to repeated training.
@@ 3 Preliminary
Continual Instruction Tuning Definition: CIT (Zhang et al., 2023) is defined to leverage LFMs to continually instruction-tune on new datasets without costly re-training. Compared with traditional continual learning, CIT differs in that it pays more attention to effectively leveraging the rich natural language instructions to prevent catastrophic forgetting and encourage knowledge transfer. Additionally, each task is an independent dataset. CIT can be described as a set of datasets $\mathcal{T}_{\text{seq}}=\{t_{1},...,t_{T}\}$ that arrives sequentially. Note that datasets in the stream can be any type and are not restricted to specific categories or domains. Each dataset $t_{j}\in\mathcal{T}_{\text{seq}}$ has a natural language instruction $I^{t_{j}}$, training set $\mathcal{D}^{t_{j}}_{\text{train}}$ and test set $\mathcal{D}^{t_{j}}_{test}$. The goal of CIT is to learn a single model $f$ from $\mathcal{T}_{\text{seq}}$ sequentially.
@@
where $\beta_{t}$ is the EMA weight, $t$ and $t-1$ is the training iteration. According to Eq.(1), performance on the current training iteration of $\theta_{t}^{*}$ is worse than $\theta_{t}$ because it only transfers a portion of the new knowledge from $\theta_{t}$ to $\theta_{t}^{*}$.
@@ 4.1 Proposition of Our Method
The first equation of Eq.(2) represents ensuring the model performance on the new dataset with no change of training loss (ideal new knowledge transferring), inspired by The Optimal Brain Surgeon framework (Hassibi et al., 1993; LeCun et al., 1989; Frantar & Alistarh, 2022; Molchanov et al., 2016). The second equation of Eq.(2) represents preserving the model performance on old datasets with no change of model parameters (ideal old knowledge protecting).
@@ 4.2 Self-Adaption Dynamical Update Method
$\text{min}\;\{\mathcal{L}(\theta_{t}^{*})-\mathcal{L}(\theta_{t})+\theta_{t}^{*}-\theta_{t-1}^{*}\},$ (7)
[…]
@@
Discussion: Based on Eq.(12), we can discover that the optimal weight is meanwhile basically related to the new gradient $\mathcal{L}^{\prime}$ and the old EMA parameters $\theta_{t-1}^{*}$. It illustrates that the obtained EMA weight $\beta_{t}$ can make the trade-off between learning new knowledge and alleviating the forgetting of old knowledge.
@@ 4.3 Two Approximate Optimizations
In Eq.(12), the calculation of ′′$\mathcal{L}^{\prime\prime}(\theta_{t})$ involves the inverse of the Hessian matrix, which needs to obtain second-order partial derivatives. However, the above calculation is complex, leading to expensive memory and time-consuming, let alone for LLM, which further increases the training burden. Thus, how to approximately express the Hessian matrix without a complex calculation process becomes a challenge and urgently needs to be solved.
Approximate Optimization Step I: Considering that the Hessian matrix is obtained by partially deriving the gradients, we can approximate the derivative with the quotient of the differentiation. Here we recognize each iteration as a period of parameter update and further simplify the denominator. As a result, we have the following approximation equation to estimate the ′′$\mathcal{L}^{\prime\prime}(\theta_{t})$.
′′$\mathcal{L}^{\prime\prime}(\theta_{t})=\frac{\partial\mathcal{L}^{\prime}(\theta_{t})}{\partial\theta_{t}}=\frac{\mathcal{L}^{\prime}(\theta_{t})-\mathcal{L}^{\prime}(\theta_{t-1})}{\theta_{t}-\theta_{t-1}},$ (13)
@@ 4.4 Instruction Grouping Strategy
Inspired by L2P and Eproj (Wang et al., 2022b; He et al., 2023)  [^3] Detailed comparisons can be found in Appendix G, we design the following strategy to decide whether to retrain or expand the training parameters. We primarily build an instruction codebook and a training pool. Instructions are grouped in the codebook and each group is associated with a set of training parameters in the pool.
@@
LoRA Fine-Tune(Hu et al., 2022) ICLR’22 21.26 28.74 10.25 36.78 32.45 0.83 42.50 57.08 28.74 37.29 61.36
MoELoRA(Chen et al., 2024a) ArXiv’24 58.92 38.59 8.85 37.10 44.25 2.45 41.40 55.35 35.86 25.71 58.36
LWF(Li & Hoiem, 2017) TPAMI’16 63.14 39.60 8.90 34.83 14.53 2.48 40.67 62.35 33.31 22.32 52.58
EWC(Kirkpatrick et al., 2017) PNAS’17 67.41 40.41 8.18 35.05 37.88 2.67 41.27 61.02 36.74 20.51 54.68
MT(Zhu et al., 2024) ICML’24 79.63 55.47 35.64 58.70 44.37 32.20 62.21 61.59 53.73 14.03 66.00
PGP(Qiao et al., 2024a) ICLR’24 85.17 56.85 32.26 61.74 49.43 32.74 65.74 62.20 55.77 12.94 67.09
EProj(He et al., 2023) ArXiv’23 78.51 57.53 92.35 55.93 44.67 36.59 63.74 57.00 60.79 5.42 65.54
@@
LoRA Fine-Tune(Hu et al., 2022) ICLR’22 60.03 41.19 10.62 31.03 32.67 2.60 46.33 61.00 35.68 32.90 64.47
MT(Zhu et al., 2024) ICML’24 80.43 60.72 46.70 60.35 49.19 33.16 63.74 65.44 57.47 11.26 67.32
PGP(Qiao et al., 2024a) ICLR’24 82.50 60.64 49.15 62.53 49.43 37.37 65.57 65.82 59.13 10.11 67.98
EProj(He et al., 2023) ArXiv’23 77.65 58.93 92.31 60.22 38.27 33.77 64.39 65.80 61.42 5.84 66.53
@@
LoRA Fine-Tune(Hu et al., 2022) ICLR’22 31.05 42.45 29.57 55.57 15.30 40.33 67.75 47.80 41.23 19.36 58.17
PGP(Qiao et al., 2024a) ICLR’24 66.42 41.33 32.16 49.83 36.05 24.22 58.60 43.96 44.07 5.90 48.30
@@ 4.5 Overview of Our Method
Our method consists of five steps, as shown in Figure 4. Before training, the initial step is to match instructions of the current task with instructions in the codebook. If we find a similar instruction group, we utilize the corresponding parameters checkpoint to initialize parameters $\theta$ and EMA parameters $\theta^{*}$. Otherwise, we randomly initialize parameters $\theta$ and EMA parameters $\theta^{*}$. After the training of iteration $t$, gradients $\mathcal{L}^{\prime}(\theta_{t})$ and model parameters $\theta_{t}$ would be saved and involved in the EMA weight calculation process of the next iteration $t+1$. Based on Eq.(14), we can obtain the adaption EMA weight $\beta_{t}$. With the EMA weight $\beta_{t}$, we can update the EMA parameters from $\theta_{t-1}^{*}$ to $\theta_{t}^{*}$ based on Eq.(1). After the training of iteration $t$, to reduce the memory burden, we will clear the saved gradients $\mathcal{L}^{\prime}(\theta_{t-1})$ and model parameters $\theta_{t-1}$. After training in each downstream dataset, we only save the EMA parameters $\theta^{*}$ and update the instruction codebook. For the detailed algorithm process please kindly refer to Appendix P.
@@ 5.1 Experimental Setup
Implementation: We adopt LLaVA-1.5/Qwen-VL (Liu et al., 2024b; qwen) as our backbone with inserted LoRA (Hu et al., 2022) in the LLM side. For all mentioned methods, including ours and other baselines, during the whole continual instruction tuning, we freeze the vision encoder and LLM, with only training the projection layer and LoRA. We follow the datasets and tuning orders of the CoIN benchmark (Chen et al., 2024a), including ScienceQA (Lu et al., 2022), TextVQA (Singh et al., 2019), ImageNet (Deng et al., 2009), GQA (Hudson & Manning, 2019), VizWiz (Gurari et al., 2018), Grounding (refcoco1; refcoco2), VQAv2 (Goyal et al., 2017) and OCR-VQA (Mishra et al., 2019). Experiments about continual instruction tuning with Large Language Models can be found in Appendix M. All experiments are conducted on 8 NVIDIA H100 GPUs.
Compared Methods: We compare our method against nine methods including (1) zero-shot and (2) LoRA Fine-Tune (Hu et al., 2022) (3) MoELoRA (Chen et al., 2024a); (4) LWF (Li & Hoiem, 2017); (5) EWC (Kirkpatrick et al., 2017); (6) EProj (He et al., 2023); (7) MT (Zhu et al., 2024); (8) PGP (Qiao et al., 2024a); (9) Multi-Task. The detailed descriptions of each method can be found in Appendix K.
Evaluation Metrics: We follow the most popular protocols for evaluation (Wang et al., 2022b, a; Smith et al., 2023; Qiao et al., 2024b), which are Average Accuracy (Simplified as Avg.Acc), Forgetting, and New Accuracy (Simplified as New.Acc). Please refer to Appendix L for more details.

### v5

@@ Abstract
Continual Instruction Tuning (CIT) is adopted to continually instruct Large Models to follow human intent data by data. It is observed that existing gradient update would heavily destroy the performance on previous datasets during CIT process. Instead, Exponential Moving Average (EMA), owns the ability to trace previous parameters, which can aid in decreasing forgetting. Nonetheless, its stable balance weight fails to deal with the ever-changing datasets, leading to the out-of-balance between plasticity and stability. In this paper, we propose a general continual instruction tuning framework to address the challenge. Starting from the trade-off prerequisite and EMA update, we propose the plasticity and stability ideal condition. Based on Taylor expansion in the loss function, we find the optimal balance weight can be automatically determined by the gradients and learned parameters. Therefore, we propose a stable-plasticity balanced coefficient to avoid knowledge interference. Based on the semantic similarity of the instructions, we can determine whether to retrain or expand the training parameters and allocate the most suitable parameters for the testing instances. Extensive experiments across multiple continual instruction tuning benchmarks demonstrate that our approach not only enhances anti-forgetting capabilities but also significantly improves overall continual tuning performance. Our code is available at https://github.com/JingyangQiao/CoIN.
@@ 1 Introduction
((a)) Radar chart of comparisons on Final Accuracy (higher is better) between baseline (LoRA Fine-Tune) and ours.
Large Foundation Modals (LFMs) have demonstrated remarkable capabilities in multi-task understanding and generation (blip2; minigpt4; gpt4; llama). It generally exits two stages: large scale pre-training and instruction-tuning. Instruction tuning is extremely significant due to guiding LFMs (instructflm; minigpt5; llama2) in following human intent and aligning different modalities (mfms; align), which seems to be an essential technique to enhance the capabilities and controllability of LFMs e.g. LLMs, MLLMs (jiang2024effectiveness).
As knowledge continuously evolves with the development of human society, new instructions are constantly generated, e.g. the emergence of new concepts or disciplines. How to enable existing LFMs to assimilate novel instructions and undergo self-evolution becomes the key challenge (mt). To accommodate the new instructions, the most effective strategy is incorporating both old and new instructions for joint training. However, even such relatively lightweight fine-tuning is unaffordable. Furthermore, directly fine-tuning these new instructions would destroy the pre-training knowledge, e.g. catastrophic forgetting (fgt1; fgt2; fgt3).
@@
Continual Instruction Tuning (CIT) is proposed to address this challenge (citb; mcit; coin). Taking MLLMs as a case study, previous methods, EProj (mcit), FwT(fwt), and CoIN (coin), utilize the model expansion framework by continually adding new branches for the novel instructions, which therefore has less impact on the old knowledge. However, they suffer from memory explosion and high computational cost problems. On the other hand, continually full fine-tuning (FFT) downstream datasets with single branch architecture would destroy the pre-trained parameters, and greatly reduce the zero-shot generalization performance of MLLMs (hal).
@@
In order to validate our method, for MLLMs, we adopt LLaVA-1.5 (llava1.5) and Qwen-VL (qwen) as the backbone and insert the efficient-tuning parameters: LoRA (lora) in the LLM. We continually fine-tune on multimodal instruction datasets and verify performances of the tuned model. Furthermore, we also implement our method with the LM-adapted version of T5-small on the NLP continual instruction tuning tasks (t5; citb). Experimental results consistently demonstrate excellent anti-forgetting, and continual tuning performance of our method. In summary, the contributions of this paper are as follows:
@@
Generalized Application and Limited Tuning Costs. Our method is model-agnostic and can be easily applied to a wide range of CIT methods. Additionally, it costs limited tuning resources, due to instruction grouping strategy.
State-of-The-Art Performance. To the best of our knowledge, our method shows the best comprehensive continual tuning performance, especially significant enhancements in the performance of the LoRA Fine-tuning baseline (as shown in Figure 3(a)).
@@ 2 Related Work
Large Foundation Models. LFMs own strong reasoning and complex contextual understanding capabilities, which can generate text sequences according to the input images and texts. With a frozen large language model and visual tower, BLIP2 (blip2) bridged the gap between image and text modality through Qformer structure. Inspired by the notable reasoning power of instruction tuning in LLMs, e.g., LLaMA (llama) and GPT (gpt3; gpt4), LLaVA (llava) and MiniGPT4 (minigpt4) adopted instruction tuning with only training a linear projection layer to align the image-text modalities. Recently, LLaVA-1.5 (llava1.5) and MiniGPT5 (minigpt5) refined the instructions and improved the performance across a wider range of multi-modality datasets.
Continual Instruction Tuning Work: (citb) first proposed the continual instruction tuning with LLMs. After that, TRACE, another continual instruction tuning benchmark, is designed to evaluate the general ability, instruction following and safety for LLMs (trace). Following them, (mcit) and (coin) designed distinct cross-modality continual instruction tuning benchmarks with InstructBLIP and LLaVA, respectively. Specifically, (mcit) adopted unique linear projection layers for each tuning dataset and utilized a key-query driven mechanism to infer the dataset identifier. (coin) employed the MOELoRA paradigm, assigning different LoRA weights to different datasets based on expert knowledge. Nevertheless, they still faced the issue of forgetting due to repeated training (yang2023continual).
@@ 3 Preliminary
Continual Instruction Tuning Definition: CIT (citb) is defined to leverage LFMs to continually instruction-tune on new datasets without costly re-training. Compared with traditional continual learning, CIT differs in that it pays more attention to effectively leveraging the rich natural language instructions to prevent catastrophic forgetting and encourage knowledge transfer. Additionally, each task is an independent dataset. CIT can be described as a set of datasets $\mathcal{T}_{\text{seq}}=\{t_{1},...,t_{T}\}$ that arrives sequentially. Note that datasets in the stream can be any type and are not restricted to specific categories or domains. Each dataset $t_{j}\in\mathcal{T}_{\text{seq}}$ has a natural language instruction $I^{t_{j}}$, training set $\mathcal{D}^{t_{j}}_{\text{train}}$ and test set $\mathcal{D}^{t_{j}}_{test}$. The goal of CIT is to learn a single model $f$ from $\mathcal{T}_{\text{seq}}$ sequentially.
@@
where $\beta_{t}$ is the EMA weight, $t$ and t1t-1 is the training iteration. According to Eq.(1), performance on the current training iteration of $\theta_{t}^{*}$ is worse than $\theta_{t}$ because it only transfers a portion of the new knowledge from $\theta_{t}$ to $\theta_{t}^{*}$.
@@ 4.1 Proposition of Our Method
The first equation of Eq.(2) represents ensuring the model performance on the new dataset with no change of training loss (ideal new knowledge transferring), inspired by The Optimal Brain Surgeon framework (obs; obd; obc; pruning). The second equation of Eq.(2) represents preserving the model performance on old datasets with no change of model parameters (ideal old knowledge protecting).
@@ 4.2 Self-Adaption Dynamical Update Method
$\text{min}\;|\mathcal{L}(\theta_{t}^{*})-\mathcal{L}(\theta_{t})+\theta_{t}^{*}-\theta_{t-1}^{*}|,$ (7)
$\textit{s.t.}\Delta\theta+\theta_{t-1}^{*}-\theta_{t}^{*}=0.$
@@
Discussion: Based on Eq.(12), we can discover that the optimal weight is basically related to the new gradient $\mathcal{L}^{\prime}$ and the old EMA parameters $\theta_{t-1}^{*}$. It illustrates that the obtained EMA weight $\beta_{t}$ can make the trade-off between learning new knowledge and alleviating the forgetting of old knowledge. For the high-dimensional matrix $\Theta_{t}$, we also have the similar formation of $\beta_{t}$ (Detailed deduction is presented in Appendix P)
@@ 4.3 Two Approximate Optimizations
In Eq.(12), the calculation of $\mathcal{L}^{\prime\prime}(\theta_{t})$ involves the inverse of the Hessian matrix, which needs to obtain second-order partial derivatives. However, the above calculation is complex, leading to expensive memory and time-consuming, let alone for LLM, which further increases the training burden. Thus, how to approximately express the Hessian matrix without a complex calculation process becomes a challenge and urgently needs to be solved.
Approximate Optimization Step I: Considering that the Hessian matrix is obtained by partially deriving the gradients, we can approximate the derivative with the quotient of the differentiation. Here we recognize each iteration as a period of parameter update and further simplify the denominator. As a result, we have the following approximation equation to estimate the $\mathcal{L}^{\prime\prime}(\theta_{t})$.
$\mathcal{L}^{\prime\prime}(\theta_{t})=\frac{\partial\mathcal{L}^{\prime}(\theta_{t})}{\partial\theta_{t}}=\frac{\mathcal{L}^{\prime}(\theta_{t})-\mathcal{L}^{\prime}(\theta_{t-1})}{\theta_{t}-\theta_{t-1}},$ (13)
@@ 4.4 Instruction Grouping Strategy
Inspired by L2P and Eproj (l2p; mcit)  [^3] Detailed comparisons can be found in Appendix G, we design the following strategy to decide whether to retrain or expand the training parameters. We primarily build an instruction codebook and a training pool. Instructions are grouped in the codebook and each group is associated with a set of training parameters in the pool.
@@
LoRA Fine-Tune(lora) ICLR’22 21.26 28.74 10.25 36.78 32.45 0.83 42.50 57.08 28.74 37.29 61.36
MoELoRA(coin) ArXiv’24 58.92 38.59 8.85 37.10 44.25 2.45 41.40 55.35 35.86 25.71 58.36
LWF(lwf) TPAMI’16 63.14 39.60 8.90 34.83 14.53 2.48 40.67 62.35 33.31 22.32 52.58
EWC(ewc) PNAS’17 67.41 40.41 8.18 35.05 37.88 2.67 41.27 61.02 36.74 20.51 54.68
MT(mt) ICML’24 79.63 55.47 35.64 58.70 44.37 32.20 62.21 61.59 53.73 14.03 66.00
PGP(pgp) ICLR’24 85.17 56.85 32.26 61.74 49.43 32.74 65.74 62.20 55.77 12.94 67.09
EProj(mcit) ArXiv’23 78.51 57.53 92.35 55.93 44.67 36.59 63.74 57.00 60.79 5.42 65.54
@@
LoRA Fine-Tune(lora) ICLR’22 60.03 41.19 10.62 31.03 32.67 2.60 46.33 61.00 35.68 32.90 64.47
MT(mt) ICML’24 80.43 60.72 46.70 60.35 49.19 33.16 63.74 65.44 57.47 11.26 67.32
PGP(pgp) ICLR’24 82.50 60.64 49.15 62.53 49.43 37.37 65.57 65.82 59.13 10.11 67.98
EProj(mcit) ArXiv’23 77.65 58.93 92.31 60.22 38.27 33.77 64.39 65.80 61.42 5.84 66.53
@@
LoRA Fine-Tune(lora) ICLR’22 31.05 42.45 29.57 55.57 15.30 40.33 67.75 47.80 41.23 19.36 58.17
PGP(pgp) ICLR’24 66.42 41.33 32.16 49.83 36.05 24.22 58.60 43.96 44.07 5.90 48.30
@@ 4.5 Overview of Our Method
Our method consists of five steps, as shown in Figure 4. Before training, the initial step is to match instructions of the current task with instructions in the codebook. If we find a similar instruction group, we utilize the corresponding parameters checkpoint to initialize parameters $\theta$ and EMA parameters $\theta^{*}$. Otherwise, we randomly initialize parameters $\theta$ and EMA parameters $\theta^{*}$. After the training of iteration $t$, gradients $\mathcal{L}^{\prime}(\theta_{t})$ and model parameters $\theta_{t}$ would be saved and involved in the EMA weight calculation process of the next iteration $t+1$. Based on Eq.(14), we can obtain the adaption EMA weight $\beta_{t}$. With the EMA weight $\beta_{t}$, we can update the EMA parameters from $\theta_{t-1}^{*}$ to $\theta_{t}^{*}$ based on Eq.(1). After the training of iteration $t$, to reduce the memory burden, we will clear the saved gradients $\mathcal{L}^{\prime}(\theta_{t-1})$ and model parameters $\theta_{t-1}$. After training in each downstream dataset, we only save the EMA parameters $\theta^{*}$ and update the instruction codebook. For the detailed algorithm process please kindly refer to Appendix Q.
@@ 5.1 Experimental Setup
Implementation: We adopt LLaVA-1.5/Qwen-VL (llava; qwen) as our backbone with inserted LoRA (lora) in the LLM side. For all mentioned methods, including ours and other baselines, during the whole continual instruction tuning, we freeze the vision encoder and LLM, with only training the projection layer and LoRA. We follow the datasets and tuning orders of the CoIN benchmark (coin), including ScienceQA (sqa), TextVQA (tqa), ImageNet (imn), GQA (gqa), VizWiz (viz), Grounding (refcoco1; refcoco2), VQAv2 (vqa) and OCR-VQA (oqa). Experiments about continual instruction tuning with Large Language Models can be found in Appendix M.
Compared Methods: We compare our method against nine methods including (1) zero-shot and (2) LoRA Fine-Tune (lora) (3) MoELoRA (coin); (4) LWF (lwf); (5) EWC (ewc); (6) EProj (mcit); (7) MT (mt); (8) PGP (pgp); (9) Multi-Task. The detailed descriptions of each method can be found in Appendix K.
Evaluation Metrics: We follow the most popular protocols for evaluation (l2p; dual; coda; pegp), which are Average Accuracy (Simplified as Avg.Acc), Forgetting, and New Accuracy (Simplified as New.Acc). Please refer to Appendix L for more details.